# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 6.2. Fourth moment bounds for $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$
(7.3) $\forall~{}\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon},\vec{\mathrm{i}}_{2}\in A_{\vec{\mathrm{i}}_{1}},\vec{\mathrm{i}}_{3}\in A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}},\quad\partial_{\vec{\mathrm{i}}_{1}\vec{\mathrm{i}}_{2}}^{2}\mathscr{L}_{\varepsilon}(\Theta)=\partial_{\vec{\mathrm{i}}_{1}\vec{\mathrm{i}}_{3}}^{2}\mathscr{L}_{\varepsilon}(\Theta)=\partial_{\vec{\mathrm{i}}_{2}\vec{\mathrm{i}}_{2}}^{3}\mathscr{L}_{\varepsilon}(\Theta)=0,$
where $\partial_{\vec{\mathrm{i}}}$ is the partial derivative with respect to $\Theta(\vec{\mathrm{i}})$. Note that $\mathscr{L}_{\varepsilon}(\Theta)$ is a polynomial of $\Theta$. We only need to verify that in each term of $\mathscr{L}_{\varepsilon}(\Theta)$, there is at most one $\Theta(\vec{\mathrm{i}}_{k})$, $k=1,2,3$. Given $\Theta(\vec{\mathrm{i}}_{1})$ and $\Theta(\vec{\mathrm{i}}_{2})$ in $\mathscr{L}_{\varepsilon}(\Theta)$, we must have $\mathrm{i}_{2}>\mathrm{i}^{\prime}_{1}$ or $\mathrm{i}_{1}>\mathrm{i}^{\prime}_{2}$. However, by the $\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}$ condition, we must have $\mathrm{i}_{2}-\mathrm{i}^{\prime}_{1}>K_{\varepsilon}$ and $\mathrm{i}_{1}-\mathrm{i}^{\prime}_{2}>K_{\varepsilon}$. Hence, $\vec{\mathrm{i}}_{2}\not\in A_{\vec{\mathrm{i}}_{1}}$, that is, no terms contain the product $\Theta(\vec{\mathrm{i}}_{1})\Theta(\vec{\mathrm{i}}_{2})$ for $\vec{\mathrm{i}}_{2}\in A_{\vec{\mathrm{i}}_{1}}$. By the same argument, given $\Theta(\vec{\mathrm{i}}_{1}),\Theta(\vec{\mathrm{i}}_{2})$ and $\Theta(\vec{\mathrm{i}}_{3})$ in $\mathscr{L}_{\varepsilon}(\Theta)$, we must have $\vec{\mathrm{i}}_{3}\notin A_{\vec{\mathrm{i}}_{1}}$ and $\vec{\mathrm{i}}_{3}\notin A_{\vec{\mathrm{i}}_{2}}$. We have now verified (7.3).
@@
(7.4) $\big{|}{\mathbbm{E}}\big{[}f(\mathscr{L}_{\varepsilon}(\Theta_{m})\big{]}-{\mathbbm{E}}\big{[}f(\mathscr{L}_{\varepsilon}(\Theta_{n}))\big{]}\big{|}\leq I_{1}^{(m)}+I_{1}^{(n)}+I_{2}^{(m)}+I_{2}^{(n)}+I_{3}^{(m,n)}.$
Here $I_{1}^{(m)},I_{1}^{(n)},I_{2}^{(m)}$ and $I_{2}^{(n)}$ comes from Lemma A.1 and $I_{3}^{(m,n)}$ come from Lemma A.2. We have applied the lemmas to $X_{k}=\Theta_{k}$, $h(\cdot)=f(\mathscr{L}_{\varepsilon}(\cdot))$ and $Z_{k}=\Theta^{\rm(G)}_{k}$ a family of centered Gaussian random variable with the same covariance structure as $\Theta_{k}$, for $k=m,n$. We will show that
(7.5) $\lim\limits_{\varepsilon\to 0}\limsup\limits_{n\to\infty}I_{1}^{(n)}$ $=0,$
(7.6) $\lim\limits_{\varepsilon\to 0}\limsup\limits_{n\to\infty}I_{2}^{(n)}$ $=0,$
(7.7) $\lim\limits_{\varepsilon\to 0}\limsup\limits_{n,m\to\infty}I_{3}^{(m,n)}$ $=0.$
@@
(7.8) $\begin{split}|I_{1}^{(n)}|\leq&\frac{1}{2}\|f^{\prime\prime\prime}\|_{\infty}\sup_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon}}{\mathbbm{E}}\big{[}|\Theta_{n}(\vec{\mathrm{i}}_{1})|^{3}\big{]}\sum\limits_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon},\vec{\mathrm{i}}_{2}\in A_{\vec{\mathrm{i}}_{1}},\vec{\mathrm{i}}_{3}\in A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}}\\ &\sup_{s,t,u}{\mathbbm{E}}\Big{[}\big{|}\partial_{\vec{\mathrm{i}}_{1}}\mathscr{L}_{\varepsilon}(W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}})\big{|}^{3}\Big{]}^{\frac{1}{3}}\sup_{s,t,u}{\mathbbm{E}}\Big{[}\big{|}\partial_{\vec{\mathrm{i}}_{2}}\mathscr{L}_{\varepsilon}(W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}})\big{|}^{3}\Big{]}^{\frac{1}{3}}\sup_{s,t,u}{\mathbbm{E}}\Big{[}\big{|}\partial_{\vec{\mathrm{i}}_{3}}\mathscr{L}_{\varepsilon}(W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}})\big{|}^{3}\Big{]}^{\frac{1}{3}},\end{split}$
where for $s,t,u\in[0,1]$,
$W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}:=su\sqrt{t}\Theta_{n}^{A_{\vec{\mathrm{i}}_{1}}}+u\sqrt{t}\Theta_{n}^{A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\backslash A_{\vec{\mathrm{i}}_{1}}}+\sqrt{t}\Theta_{n}^{A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}^{c}}+\sqrt{1-t}\Theta_{n}^{\rm(G)},\quad\text{with}~{}\Theta_{n}^{A}(\vec{\mathrm{i}}):=\Theta_{n}(\vec{\mathrm{i}})\mathbbm{1}_{\{\vec{\mathrm{i}}\in A\}}.$
First note that there are finite many terms in $\mathscr{L}_{\varepsilon}(\Theta_{n})$, since $\varphi,\psi\in C_{c}({\mathbbm{R}})$. Hence, the number of indices $\vec{\mathrm{i}}$ is finite and we can pass the $\lim_{n\to\infty}$ inside the summation. By Lemma 6.2 and the boundedness of $f^{\prime\prime\prime}$, we readily have that
(7.9) $\limsup_{n\to\infty}\sup_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon}}\|f^{\prime\prime\prime}\|_{\infty}{\mathbbm{E}}\big{[}|\Theta_{n}(\vec{\mathrm{i}}_{1})|^{3}\big{]}\leq C\limsup_{n\to\infty}\sup_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon}}{\mathbbm{E}}\big{[}|\Theta_{n}(\vec{\mathrm{i}}_{1})|^{4}\big{]}^{\frac{3}{4}}\leq\frac{C}{(\log\frac{1}{\varepsilon})^{\frac{3}{4}}}.$
@@
(7.10) $\sum\limits_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon},\vec{\mathrm{i}}_{2}\in A_{\vec{\mathrm{i}}_{1}},\vec{\mathrm{i}}_{3}\in A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}}\frac{1}{3}\sum\limits_{k=1}^{3}\sup_{s,t,u}{\mathbbm{E}}\Big{[}\Big{|}\partial_{\vec{\mathrm{i}}_{k}}\mathscr{L}_{\varepsilon}\Big{(}W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big{)}\Big{|}^{4}\Big{]}^{\frac{3}{4}}.$
@@
(7.11) $W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}(\vec{\mathrm{i}})=\begin{cases}su\sqrt{t}\Theta_{n}(\vec{\mathrm{i}})+\sqrt{1-t}\Theta_{n}^{\rm(G)}(\vec{\mathrm{i}}),\quad&\text{for}~{}\vec{\mathrm{i}}\in A_{\vec{\mathrm{i}}_{1}},\\ u\sqrt{t}\Theta_{n}(\vec{\vec{\mathrm{i}}})+\sqrt{1-t}\Theta_{n}^{\rm(G)}(\vec{\mathrm{i}}),\quad&\text{for}~{}\vec{\mathrm{i}}\in A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\backslash A_{\vec{\mathrm{i}}_{1}},\\ \sqrt{t}\Theta_{n}(\vec{\mathrm{i}})+\sqrt{1-t}\Theta_{n}^{\rm(G)}(\vec{\mathrm{i}}),\quad&\text{for}~{}\vec{\mathrm{i}}\in A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}^{c}.\end{cases}$
@@
$\begin{split}\mathscr{L}_{\varepsilon}\Big{(}W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}\Big{)}=g_{1}(\varphi,\psi)+&\sqrt{\varepsilon}\sum\limits_{r=1}^{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}}\\ &\varphi_{\varepsilon}(\mathrm{a})g_{\mathrm{i}_{1}}(\mathrm{a})W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}}_{1})\times\bigg{\{}\prod_{j=2}^{r}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)W_{s,t,u}^{\vec{\mathrm{j}}_{i},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}}_{j})\bigg{\}}g_{(\lfloor\frac{1}{\varepsilon}\rfloor-\mathrm{i}^{\prime}_{r})}(\mathrm{b})\psi_{\varepsilon}(\mathrm{b}),\end{split}$
@@
(7.12) $\partial_{\vec{\mathrm{i}}}W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}})=\begin{cases}su\sqrt{t},\quad&\text{for}~{}\vec{\mathrm{i}}\in A_{\vec{\mathrm{j}}_{1}},\\ u\sqrt{t},\quad&\text{for}~{}\vec{\mathrm{i}}\in A_{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}\backslash A_{\vec{\mathrm{j}}_{1}},\\ \sqrt{t},\quad&\text{for}~{}\vec{\mathrm{i}}\in A_{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}^{c}.\end{cases}$
@@
(7.13) $\partial_{\vec{\mathrm{i}}}\mathscr{L}_{\varepsilon}\Big{(}W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big{)}=C(\vec{\mathrm{i}})\frac{1}{\sqrt{\varepsilon}}\overline{\mathscr{L}}_{\varepsilon,[0,\mathrm{i}]}\Big{(}\varphi,\delta_{0}|W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big{)}\overline{\mathscr{L}}_{\varepsilon,[\mathrm{i}^{\prime},\frac{1}{\varepsilon}]}\Big{(}\delta_{0},\psi|W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big{)},$
where $C(\mathrm{i})$ is the constant determined by (7.12), $\delta_{0}$ is the Dirac measure, and ¯$\overline{\mathscr{L}}_{\varepsilon,[t_{1},t_{2}]}$ is a coarse-grained model without the deterministic term and having the same structure as $\mathscr{L}_{\varepsilon}$ with initial time $t_{1}$ and terminal time $t_{2}$. Here we need to multiply an $1/\sqrt{\varepsilon}$ because we decompose one coarse-grained model by two coarse-grained models, and thus an extra $\sqrt{\varepsilon}$ is needed by the definition of $\mathscr{L}_{\varepsilon}$.
@@
(7.14) ${\mathbbm{E}}\Big{[}\Big{|}\partial_{\vec{\mathrm{i}}}\mathscr{L}_{\varepsilon}\Big{(}W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big{)}\Big{|}^{4}\Big{]}\leq\frac{1}{\varepsilon^{2}}{\mathbbm{E}}\Big{[}\Big{|}\overline{\mathscr{L}}_{\varepsilon,[0,\mathrm{i}]}\Big{(}\delta_{0},\varphi|W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big{)}\Big{|}^{4}\Big{]}{\mathbbm{E}}\Big{[}\Big{|}\overline{\mathscr{L}}_{\varepsilon,[\mathrm{i}^{\prime},\frac{1}{\varepsilon}]}\Big{(}\delta_{0},\psi|W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big{)}\Big{|}^{4}\Big{]},$
where we have used $C(\mathrm{i})\leq 1$ by $s,u,t\in[0,1]$ and the time reversal invariance for ¯$\overline{\mathscr{L}}_{\varepsilon}$ for the first expectation. Note that ¯$\overline{\mathscr{L}}_{\varepsilon}(W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}})$ has mean 00 and we have the following analogue of Theorem 6.3 for coarse-grained disorders $W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}$.

### v3

@@ 6.1 Moment bounds for the coarse-grained variable $\Theta_{N,\varepsilon}$
$\forall~\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon},\vec{\mathrm{i}}_{2}\in A_{\vec{\mathrm{i}}_{1}},\vec{\mathrm{i}}_{3}\in A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}},\quad\partial_{\vec{\mathrm{i}}_{1}\vec{\mathrm{i}}_{2}}^{2}\mathscr{L}_{\varepsilon}(\Theta)=\partial_{\vec{\mathrm{i}}_{1}\vec{\mathrm{i}}_{3}}^{2}\mathscr{L}_{\varepsilon}(\Theta)=\partial_{\vec{\mathrm{i}}_{2}\vec{\mathrm{i}}_{3}}^{2}\mathscr{L}_{\varepsilon}(\Theta)=0,$ (7.3)
where $\partial_{\vec{\mathrm{i}}}$ is the partial derivative with respect to $\Theta(\vec{\mathrm{i}})$. Note that $\mathscr{L}_{\varepsilon}(\Theta)$ is a polynomial of $\Theta$. We only need to verify that in each term of $\mathscr{L}_{\varepsilon}(\Theta)$, there is at most one $\Theta(\vec{\mathrm{i}}_{k})$, $k=1,2,3$. Given $\Theta(\vec{\mathrm{i}}_{1})$ and $\Theta(\vec{\mathrm{i}}_{2})$ in $\mathscr{L}_{\varepsilon}(\Theta)$, we must have $\mathrm{i}_{2}>\mathrm{i}^{\prime}_{1}$ or $\mathrm{i}_{1}>\mathrm{i}^{\prime}_{2}$. However, by the $\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}$ condition, we must have $\mathrm{i}_{2}-\mathrm{i}^{\prime}_{1}>K_{\varepsilon}$ or $\mathrm{i}_{1}-\mathrm{i}^{\prime}_{2}>K_{\varepsilon}$. Hence, $\vec{\mathrm{i}}_{2}\not\in A_{\vec{\mathrm{i}}_{1}}$, that is, no terms contain the product $\Theta(\vec{\mathrm{i}}_{1})\Theta(\vec{\mathrm{i}}_{2})$ for $\vec{\mathrm{i}}_{2}\in A_{\vec{\mathrm{i}}_{1}}$. By the same argument, given $\Theta(\vec{\mathrm{i}}_{1}),\Theta(\vec{\mathrm{i}}_{2})$ and $\Theta(\vec{\mathrm{i}}_{3})$ in $\mathscr{L}_{\varepsilon}(\Theta)$, we must have $\vec{\mathrm{i}}_{3}\notin A_{\vec{\mathrm{i}}_{1}}$ and $\vec{\mathrm{i}}_{3}\notin A_{\vec{\mathrm{i}}_{2}}$. We have now verified (7.3).
@@
$\big|{\mathbbm{E}}\big[f(\mathscr{L}_{\varepsilon}(\Theta_{m})\big]-{\mathbbm{E}}\big[f(\mathscr{L}_{\varepsilon}(\Theta_{n}))\big]\big|\leq I_{1}^{(m)}+I_{1}^{(n)}+I_{2}^{(m)}+I_{2}^{(n)}+I_{3}^{(m,n)}.$ (7.4)
Here $I_{1}^{(m)},I_{1}^{(n)},I_{2}^{(m)}$ and $I_{2}^{(n)}$ comes from Lemma A.1 and $I_{3}^{(m,n)}$ come from Lemma A.2. We have applied the lemmas to $X_{k}=\Theta_{k}$, $h(\cdot)=f(\mathscr{L}_{\varepsilon}(\cdot))$ and $Z_{k}=\Theta^{\rm(G)}_{k}$, where $Z_{k}$ is a family of centered Gaussian random variables with the same covariance structure as $\Theta_{k}$, for $k=m,n$. We will show that
$\lim\limits_{\varepsilon\to 0}\limsup\limits_{n\to\infty}I_{1}^{(n)}$ $=0,$ (7.5)
$\lim\limits_{\varepsilon\to 0}\limsup\limits_{n\to\infty}I_{2}^{(n)}$ $=0,$ (7.6)
$\lim\limits_{\varepsilon\to 0}\limsup\limits_{n,m\to\infty}I_{3}^{(m,n)}$ $=0.$ (7.7)
@@
$\begin{split}|I_{1}^{(n)}|\leq&\frac{1}{2}\|f^{\prime\prime\prime}\|_{\infty}\sup_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon}}{\mathbbm{E}}\big[|\Theta_{n}(\vec{\mathrm{i}}_{1})|^{3}\big]\sum\limits_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon},\vec{\mathrm{i}}_{2}\in A_{\vec{\mathrm{i}}_{1}},\vec{\mathrm{i}}_{3}\in A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}}\\ &\sup_{s,t,u}{\mathbbm{E}}\Big[\big|\partial_{\vec{\mathrm{i}}_{1}}\mathscr{L}_{\varepsilon}(W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}})\big|^{3}\Big]^{\frac{1}{3}}\sup_{s,t,u}{\mathbbm{E}}\Big[\big|\partial_{\vec{\mathrm{i}}_{2}}\mathscr{L}_{\varepsilon}(W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}})\big|^{3}\Big]^{\frac{1}{3}}\sup_{s,t,u}{\mathbbm{E}}\Big[\big|\partial_{\vec{\mathrm{i}}_{3}}\mathscr{L}_{\varepsilon}(W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}})\big|^{3}\Big]^{\frac{1}{3}},\end{split}$ (7.8)
where for $s,t,u\in[0,1]$ and $\Theta_{n}^{A}(\vec{\mathrm{i}}):=\Theta_{n}(\vec{\mathrm{i}})\mathbbm{1}_{\{\vec{\mathrm{i}}\in A\}}$,
[…]
First note that there are finite many terms in $\mathscr{L}_{\varepsilon}(\Theta_{n})$, since $\varphi,\psi\in C_{c}({\mathbbm{R}})$. Hence, the number of indices $\vec{\mathrm{i}}$ is finite and we can pass $\lim_{n\to\infty}$ inside the summation. By Lemma 6.2 and the boundedness of $f^{\prime\prime\prime}$, we readily have that
$\limsup_{n\to\infty}\sup_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon}}\|f^{\prime\prime\prime}\|_{\infty}{\mathbbm{E}}\big[|\Theta_{n}(\vec{\mathrm{i}}_{1})|^{3}\big]\leq C\limsup_{n\to\infty}\sup_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon}}{\mathbbm{E}}\big[|\Theta_{n}(\vec{\mathrm{i}}_{1})|^{4}\big]^{\frac{3}{4}}\leq\frac{C}{(\log\frac{1}{\varepsilon})^{\frac{3}{4}}}.$ (7.9)
@@
$\sum\limits_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon},\vec{\mathrm{i}}_{2}\in A_{\vec{\mathrm{i}}_{1}},\vec{\mathrm{i}}_{3}\in A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}}\frac{1}{3}\sum\limits_{k=1}^{3}\sup_{s,t,u}{\mathbbm{E}}\Big[\Big|\partial_{\vec{\mathrm{i}}_{k}}\mathscr{L}_{\varepsilon}\Big(W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big)\Big|^{4}\Big]^{\frac{3}{4}}.$ (7.10)
@@
[…]
@@
$\begin{split}\mathscr{L}_{\varepsilon}\Big(W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}&\Big)=g_{1}(\varphi,\psi)+\sqrt{\varepsilon}\sum\limits_{r=1}^{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}}\\ &\varphi_{\varepsilon}(\mathrm{a})g_{\mathrm{i}_{1}}(\mathrm{a})W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}}_{1})\times\bigg\{\prod_{j=2}^{r}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)W_{s,t,u}^{\vec{\mathrm{j}}_{i},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}}_{j})\bigg\}g_{(\lfloor\frac{1}{\varepsilon}\rfloor-\mathrm{i}^{\prime}_{r})}(\mathrm{b})\psi_{\varepsilon}(\mathrm{b}),\end{split}$
@@
[…]
@@
$\partial_{\vec{\mathrm{i}}}\mathscr{L}_{\varepsilon}\Big(W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big)=C(\vec{\mathrm{i}})\frac{1}{\sqrt{\varepsilon}}\overline{\mathscr{L}}_{\varepsilon,[0,\mathrm{i}]}\Big(\varphi,\delta_{0}\Big|W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big)\overline{\mathscr{L}}_{\varepsilon,[\mathrm{i}^{\prime},\frac{1}{\varepsilon}]}\Big(\delta_{0},\psi\Big|W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big),$ (7.13)
where $C(\mathrm{i})$ is the constant determined by (7.12), $\delta_{0}$ is the Dirac measure, and $\overline{\mathscr{L}}_{\varepsilon,[t_{1},t_{2}]}$ is a coarse-grained model without the deterministic term and having the same structure as $\mathscr{L}_{\varepsilon}$ with initial time $t_{1}$ and terminal time $t_{2}$. Here we need to multiply an $1/\sqrt{\varepsilon}$ because we decompose one coarse-grained model by two coarse-grained models, and thus an extra $\sqrt{\varepsilon}$ is needed by the definition of $\mathscr{L}_{\varepsilon}$.
@@
${\mathbbm{E}}\Big[\Big|\partial_{\vec{\mathrm{i}}}\mathscr{L}_{\varepsilon}\Big(W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big)\Big|^{4}\Big]\leq\frac{1}{\varepsilon^{2}}{\mathbbm{E}}\Big[\Big|\overline{\mathscr{L}}_{\varepsilon,[0,\mathrm{i}]}\Big(\delta_{0},\varphi|W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big)\Big|^{4}\Big]{\mathbbm{E}}\Big[\Big|\overline{\mathscr{L}}_{\varepsilon,[\mathrm{i}^{\prime},\frac{1}{\varepsilon}]}\Big(\delta_{0},\psi|W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big)\Big|^{4}\Big],$ (7.14)
where we have used $C(\mathrm{i})\leq 1$ by $s,u,t\in[0,1]$ and the time reversal invariance for $\overline{\mathscr{L}}_{\varepsilon}$ for the first expectation. Note that $\overline{\mathscr{L}}_{\varepsilon}(W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}})$ has mean 0 and we have the following analogue of Theorem 6.3 for coarse-grained disorders $W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}$.